40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 13 21:30:55 2020, shruti, Update, ALS, ALS OOL noise with arms locked 20200213_ALS.pdf
    Reply  Fri Feb 14 14:02:13 2020, shruti, Update, ALS, ALS OOL noise with arms locked 819472_10.pdf225642_10.pdf693500_10.pdf
       Reply  Fri Feb 14 14:52:41 2020, gautam, Update, ALS, ALS OOL noise with arms locked 
          Reply  Tue Feb 18 18:14:59 2020, shruti, Update, ALS, ALS OOL noise with arms locked ALSNoise.pdf
             Reply  Wed Feb 19 22:20:22 2020, rana, Update, ALS, ALS OOL noise with arms locked 
             Reply  Fri Feb 21 10:59:08 2020, shruti, Update, ALS, PDH error signals? IMG_0082.jpgIMG_0083.jpgIMG_0084.jpgIMG_0085.jpg
                Reply  Fri Feb 21 13:02:53 2020, Koji, Update, ALS, PDH error signals? 
                   Reply  Fri Feb 21 20:44:18 2020, shruti, Update, ALS, ALS OOL noise and PDH ALSNoise.pdfIMG_0086.jpg
                      Reply  Sun Feb 23 18:15:22 2020, rana, Update, ALS, ALS OOL noise and PDH 
Message ID: 15220     Entry time: Fri Feb 21 20:44:18 2020     In reply to: 15219     Reply to this: 15221
Author: shruti 
Type: Update 
Category: ALS 
Subject: ALS OOL noise and PDH 

[Meenakshi, Shruti]

In order to adjust the relative phase for PDH locking, we used the Siglent SDG 1032X function generator which has two outputs whose relative phase can be adjusted.

This Siglent function generator was borrowed from Yehonathan's setup near the PSL table and can be found at the X end disconnected from our setup after our use.

Initially, we used the Siglent at 231.250 kHz and 5 Vpp from each output with zero relative phase to lock the green arm cavity. By moving the phase at intervals of 5deg and looking at the PDH error signals when the cavity was unlocked we concluded that 0deg probably looked like it had the largest linear region (~1.9 V on the yaxis. Refer elog 15218 for more information) as expected.

Then we tried the same for 225.642 kHz, 5 Vpp, and found the optimal demod phase to be -55deg, with linear region of ~3 V (Ref. Attachment 2). A 'bad' frequency 180 kHz was optimized to 10deg and linear region of ~1.5 V.

The error signals at higher frequencies appeared to be quite low (not sure why at the moment) and tuning the phase did not seem to help this much.

For the noise measurement, the IFO arms were locked to IR and green, but even after optimizing the transmission with dither, we couldn't achieve best locking (green transmission was around ~0.2). Further, the IMC went out of lock during the experiment after which Koji helped us by adjusting the gains a locking point of the PMC servo. Attachment 1 contains some noise curves for the 3 frequencies with a reference from an earlier 'good' time.

Attachment 1: ALSNoise.pdf  307 kB  | Hide | Hide all
ALSNoise.pdf
Attachment 2: IMG_0086.jpg  377 kB  | Hide | Hide all
IMG_0086.jpg
ELOG V3.1.3-